OpenStudy (anonymous):

2x^3+22x^2+36x solve the cubic equation by factoring out the gcf first

OpenStudy (unklerhaukus):

gcf is 2x

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

what do you have when you have factored it out

OpenStudy (anonymous):

i relly dont know

OpenStudy (anonymous):

2x(x^2+11x+18) 2x(x-9)(x-2) this is all he wants @UnkleRhaukus

OpenStudy (unklerhaukus):

\[2x^3+22x^2+36x\Longrightarrow 2x\left(\frac{2x^3}{2x}+\frac{22x^2}{2x}+\frac{36x}{2x}\right)\]

OpenStudy (unklerhaukus):

yeah but thats hardly an explanation theyatin

OpenStudy (anonymous):

please ellaborate

OpenStudy (unklerhaukus):

cancel the gfc like this \[2x\left(\frac{2x^3}{2x}+\frac{22x^2}{2x}+\frac{36x}{2x}\right)\Longrightarrow 2x\left(\frac{\cancel {2x}\times x^{2}}{\cancel{2x}}+\frac{\cancel{2x}\times11x}{\cancel{2x}}+\frac{\cancel{2x}\times16}{\cancel{2x}}\right)\]

OpenStudy (unklerhaukus):

whops 36≠2x16 36=2x18

OpenStudy (unklerhaukus):

so now you have a quadratic in the brackets

OpenStudy (anonymous):

k

OpenStudy (unklerhaukus):

the solutions for x are the values x can be for the equation to be true , but you dont have an equation

OpenStudy (anonymous):

okthnk

OpenStudy (unklerhaukus):

=0?